# Decomposing Global Feature Effects Based on Feature Interactions

### v2

@@ 4.4 GADGET-ALE
The local feature effects of GADGET-ALE satisfy the local decomposibility axiom (Axiom 1). The proof is given in Appendix A.5.
The derivatives in ALE are typically estimated by defining $m$ intervals for feature $\mathbf{x}_{j}$ and quantifying for each interval the prediction difference between the upper and lower boundary for observations lying in this interval (see Eq. 4). Hence, the loss function in Eq. (6) measures the variance of the derivatives of observations where the feature values of $\mathbf{x}_{j}$ lie within the boundaries of the regarded interval and, thus, measures the interaction-related heterogeneity of local effects of feature $\mathbf{x}_{j}$ within this interval. The risk function in Eq. (7) then aggregates this loss over all $m$ intervals.
Equivalently to PD plots, ALE plots can be decomposed additively into main and interaction effects. Furthermore, if all feature interactions of features in $S$ can be completely minimized (see Theorem 3), the joint mean-centered ALE function $f^{ALE,c}_{S|\mathcal{A}_{g}}$ within each final subspace $\mathcal{A}_{g}$ can be decomposed into the 1-dimensional mean-centered ALE functions of features in $S$ (see Eq. 8). More details on the decomposition including an illustrative example can be found in Appendix B.2.
@@ 4.5 GADGET-SD
The Shapley values satisfy the local decomposibility axiom (Axiom 1). The proof can be found in Appendix A.6.
It is important to note that the Shapley value is defined via expected values and thus its definition differs from those of ICE curves and derivatives for ALE, which are based on local predictions. Similarly to estimating the mean-centering constants for ICE curves, it follows that the expected values within the Shapley value estimation must consider the regarded subspace $\mathcal{A}_{g}$ to acknowledge the full heterogeneity reduction due to interactions within each subspace. This means that we must recalculate the Shapley values after each partitioning step for each new subspace to receive regional SD effects in the final subspaces that are representative of the underlying main effects within each subspace. However, without recalculating the conditional expected values, we still minimize the unconditional expected value (i.e., the feature interactions on the entire feature space). The two different approaches lead to a different acknowledgment of interaction effects. In general, it can be said that the faster approach without recalculation will be less likely to detect feature interactions of higher order compared to the exact approach with recalculation. The differences of the two approaches are explained in Appendix B.3 and further discussed on a more complex simulation example in Section 6.2.
@@
The decomposition of Shapley values differs from that of mean-centered ICE curves in the way that feature interactions in Shapley values are fairly distributed between involved features, which leads to decreasing weights the higher the order of the interaction effect (see Eq. 10). In contrast, all interaction terms receive the same weight as the main effects in ICE curves. Hence, interactions of high order lead to less heterogeneity in SD plots compared to ICE plots. However, by using the marginal-based approach to calculate Shapley values, they can be decomposed by weighted PD functions (see Eq. 10 and Herren and Hahn, 2022). Hence, the same decomposition rules as defined for PD plots apply for the SD feature effect, as defined in Eq. (11). Meaning, the regional joint SD effect of features in $S$ can be decomposed into 1-dimensional regional SD effect functions, as in Eq. (8) if all interaction-related heterogeneity is reduced. In this special case and if Shapley values are estimated by the marginal-based approach, it can be shown that $f^{SD,c}_{j|\mathcal{A}_{g}}=f^{PD,c}_{j|\mathcal{A}_{g}}$ (see Appendix B.3 for more details and an illustration in a simulated example).
@@ 4.6 Illustrative Comparison of Feature Effect Methods Within GADGET
Definitions and descriptions of the mean and interaction-related heterogeneity estimates and visualization techniques for each feature effect method can be found in Appendix B.4.
@@ 4.7 Quantifying Feature Interactions
$I(\mathcal{A}_{P},\tilde{\mathbf{x}}_{j})=\frac{\mathcal{R}(\mathcal{A}_{P},\tilde{\mathbf{x}}_{j})-\mathcal{R}(\mathcal{A}_{l},\tilde{\mathbf{x}}_{j})-\mathcal{R}(\mathcal{A}_{r},\tilde{\mathbf{x}}_{j})}{\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})},$ (12)
@@
$I_{z}=\frac{\sum_{\mathcal{A}_{P}\in\mathcal{B}_{z}}\sum_{j\in S}(\mathcal{R}(\mathcal{A}_{P},\tilde{\mathbf{x}}_{j})-\mathcal{R}(\mathcal{A}_{l},\tilde{\mathbf{x}}_{j})-\mathcal{R}(\mathcal{A}_{r},\tilde{\mathbf{x}}_{j}))}{\sum_{j\in S}\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})}.$ (14)
@@
$R^{2}_{j}=\sum_{z\in Z}I_{z,j}(\tilde{\mathbf{x}}_{j})=1-\frac{\sum_{\mathcal{A}_{t}\in\mathcal{B}_{t}}\mathcal{R}(\mathcal{A}_{t},\tilde{\mathbf{x}}_{j})}{\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})},$ (15)
@@
$R^{2}_{Tot}=\sum_{z\in Z}I_{z}=1-\frac{\sum_{\mathcal{A}_{t}\in\mathcal{B}_{t}}\sum_{j\in S}\mathcal{R}(\mathcal{A}_{t},\tilde{\mathbf{x}}_{j})}{\sum_{j\in S}\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})}.$ (16)
@@
5 A Significance Test for Global Feature Interactions
@@
As shown in the previous section, GADGET quantifies feature interactions between the feature subsets $S$ and $Z$. Choosing the features to include in $S$ and $Z$ depends on the underlying research question. If the user is interested in how a specific set of features ($S$) influences the model’s predictions depending on another user-defined feature set ($Z$), then $S$ and $Z$ are chosen based on domain knowledge. However, the user may not know which feature effects and interactions the ML model has learned. Thus, choosing $S$ and $Z$ based on domain knowledge does not ensure consideration of all interacting features. Thus, if our goal is to minimize feature interactions between all features to potentially additively decompose the prediction function into univariate effects, we can define $S=Z=\{1,\ldots,p\}$. With that choice, we aim to reduce the overall interaction-related heterogeneity in all features ($S$), since we also consider all features to be possible interacting features ($Z$). For high-dimensional settings, this has the disadvantage of being computationally expensive and we might “detect” spurious interactions and therefore produce less stable regional splits. Hence, we must define beforehand the subset of features that actually interact. Given that features typically interact with each other and all involved features will usually show heterogeneity in their local effects while being responsible for the heterogeneity of other involved features, we will usually choose $S=Z$. Thus, we will only use $S$ as the globally interacting feature subset to be defined.
5.1 The PINT Test
We introduce a new statistical permutation interaction (PINT) test to define the interacting feature subset $S$ based on a non-parametric permutation test (Ernst, 2004). It can be applied to any feature effect method satisfying Axiom 1. More formally, to detect if $\mathbf{x}_{j}$ is interacting with any features in $\mathbf{x}_{-j}$, we test the null hypothesis that no feature interactions are present in the true underlying function between $\mathbf{x}_{j}$ and $\mathbf{x}_{-j}$. Since the feature interaction between $\mathbf{x}_{j}$ and $\mathbf{x}_{-j}$ on a global level can be measured by the risk function $\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})$ on the entire feature space $\mathcal{X}$ (see Eq. (7)), the null hypothesis can be defined by $H_{0}:\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})=0$. If the interaction-related heterogeneity of $\mathbf{x}_{j}$ in the true functional relationship of the data-generating process is 00, then no interactions are present, and $\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})=0$ holds. As the risk function is defined by a sum of variances, its value is always larger or equal to 00. Thus, the alternative hypothesis is $H_{1}:\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})>0$, which leads to a one-sided test at significance level $\alpha$.
@@ 5.1 The PINT Test
Generally, we do not know the true functional relationship nor the distribution of the risk values under the null hypothesis. Therefore, we use the prediction model and a permutation-based approach to approximate the null distribution. We follow a similar approach to Altmann et al. (2010) who suggested a permutation-based null distribution to detect important features. To obtain the null distribution, the target variable $y$ is permuted (line 4 in Algorithm 2), which breaks the association between the target and the features while keeping the feature distribution intact. Hence, no more interactions are present between $\mathbf{x}_{j}$ and $\mathbf{x}_{-j}$, but the association between the features remains. Thus, when performing a refit on the data set $\tilde{\mathcal{D}}$ with the permuted target variable, potential heterogeneity of local effects is only due to randomly learned effects of the prediction model depending on the feature distribution and noise and not caused by true feature interactions. Therefore, calculating ~$\tilde{\mathcal{R}}_{j}$ based on $\tilde{\mathcal{D}}$ measures the risk of $\mathbf{x}_{j}$ under the null (lines 5-8 in Algorithm 2). To obtain the null distribution of risk values for $\mathbf{x}_{j}$, this procedure is repeated $s$ times. [^7] An alternative approach would be to randomly permute $\mathbf{x}_{j}$ instead of $y$ to break the interaction effects between $\mathbf{x}_{j}$ and $\mathbf{x}_{-j}$. However, this approach is computationally more expensive and does not only break the respective interaction effects but also the correlations between $\mathbf{x}_{j}$ and $\mathbf{x}_{-j}$, and thus spurious interactions will not be detected. Using conditional samples (Molnar et al., 2023) or knock-offs (Watson and Wright, 2021) instead of random permutations solves the second problem. However, this approach is even more computationally expensive and becomes infeasible with a higher number of features.
@@
Since we are interested in defining the feature subset $S$ of all interacting features, we calculate ~$\tilde{\mathcal{R}}_{j}$ for all features $j\in\{1,\ldots,p\}$. Since all feature interactions between $y$ and $\mathbf{x}$ are broken by permuting $y$, the null distribution of risk values can be estimated for all features by the described $s$ permuting and refitting steps. To determine the interacting feature subset $S$, the test statistic for each feature, so the respective risk value $\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})$, is calculated based on the unpermuted data set $\mathcal{D}$ and the originally fitted model on this data set. The test statistic value $\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})$ is then compared to the $(1-\alpha)$-quantile, i.e., $z^{1-\alpha}_{j}$, of the non-parametric null distribution. If the test statistic value of $\mathbf{x}_{j}$ exceeds the respective quantile, then $H_{0}$ is rejected and feature interactions between $\mathbf{x}_{j}$ and $\mathbf{x}_{-j}$ are considered to be significant at the significance level $\alpha$ (lines 11-13 in Algorithm 2).
@@
risk function $\mathcal{R}$, significance level $\alpha$
@@
5:     refit model on $\tilde{\mathcal{D}}^{k}$ to obtain the prediction function $\tilde{f}^{k}$
@@
7:        calculate risk ~$\tilde{\mathcal{R}}_{j}^{k}=\tilde{\mathcal{R}}^{k}(\mathcal{X},\tilde{\mathbf{x}}_{j})$ for $j$-the feature based on $\tilde{\mathcal{D}}^{k}$ and $\tilde{f}^{k}$
@@
11:     calculate risk $\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})$ for $j$-the feature based on $\mathcal{D}$ and $\hat{f}$
@@
13:     add $j$ to $S$, if and only if $\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})>z^{1-\alpha}_{j}$
@@
5.2 Theoretical Validity
A difficulty in formally analyzing the test is that it depends on the unspecified ML algorithm that induces $\hat{f}$. We can emphasize this by writing $\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}_{n}})$ for the risk for feature value $\mathbf{x}_{j}$ achieved by the algorithm $\hat{f}$ run on a data set [^8] With a slight abuse of notation, we use $\mathcal{D}_{n}$ to define the dataset as a random variable. $\mathcal{D}_{n}=\{(X^{(i)},Y^{(i)})\}_{i=1}^{n}$, generated i.i.d. from $\mathds{P}_{X,Y}$. Define $f(\mathbf{x})=\mathds{E}[Y\mid X=\mathbf{x}]$ and $\mathcal{D}^{\Pi}_{n}=\{(X^{(i)},Y_{\Pi(i)}^{(i)})\}_{i=1}^{n}$ where $\Pi$ is a random permutation of $\{1,\dots,n\}$ independent of $\mathcal{D}_{n}$. We can now impose abstract, but interpretable conditions on the ML algorithm under which the test is valid.

### v3

@@ 4.4 GADGET-ALE
The local feature effects of GADGET-ALE satisfy the local decomposability axiom (Axiom 1). The proof is given in Appendix B.5.
The derivatives in ALE are typically estimated by defining $m$ intervals for feature $\mathbf{x}_{j}$ and quantifying for each interval the prediction difference between the upper and lower boundary for observations lying in this interval (see Eq. 4). Hence, the loss function in Eq. (6) measures the variance of the derivatives of observations where the feature values of $\mathbf{x}_{j}$ lie within the boundaries of the regarded interval and, thus, measures the interaction-related heterogeneity of local effects of feature $\mathbf{x}_{j}$ within this interval. The risk function in Eq. (7) then aggregates this loss over all relevant intervals.
Equivalently to PD plots, ALE plots can be decomposed additively into main and interaction effects. Furthermore, if all feature interactions of features in $S$ can be completely minimized (see Theorem 3), the joint mean-centered ALE function $f^{ALE,c}_{S|\mathcal{A}_{g}}$ within each final subspace $\mathcal{A}_{g}$ can be decomposed into the 1-dimensional mean-centered ALE functions of features in $S$ (see Eq. 8). More details on the decomposition including an illustrative example can be found in Appendix C.2.
@@ 4.5 GADGET-SD
The Shapley values satisfy the local decomposability axiom (Axiom 1). The proof can be found in Appendix B.6.
It is important to note that the Shapley value is defined via expected values and thus its definition differs from those of ICE curves and derivatives for ALE, which are based on local predictions. Similarly to estimating the mean-centering constants for ICE curves, it follows that the expected values within the Shapley value estimation must consider the regarded subspace $\mathcal{A}_{g}$ to acknowledge the full heterogeneity reduction due to interactions within each subspace. This means that we must recalculate the Shapley values after each partitioning step for each new subspace to receive regional SD effects in the final subspaces that are representative of the underlying main effects within each subspace. However, without recalculating the conditional expected values, we still minimize the unconditional expected value (i.e., the feature interactions on the entire feature space). The two different approaches lead to a different acknowledgment of interaction effects. In general, it can be said that the faster approach without recalculation will be less likely to detect feature interactions of higher order compared to the exact approach with recalculation. The differences of the two approaches are explained in Appendix C.3 and further discussed on a more complex simulation example in Section 6.2.
@@
The decomposition of Shapley values differs from that of mean-centered ICE curves in the way that feature interactions in Shapley values are fairly distributed between involved features, which leads to decreasing weights the higher the order of the interaction effect (see Eq. 10). In contrast, all interaction terms receive the same weight as the main effects in ICE curves. Hence, interactions of high order lead to less heterogeneity in SD plots compared to ICE plots. However, by using the marginal-based approach to calculate Shapley values, they can be decomposed by weighted PD functions (see Eq. 10 and Herren and Hahn, 2022). Hence, the same decomposition rules as defined for PD plots apply for the SD feature effect, as defined in Eq. (11). Meaning, the regional joint SD effect of features in $S$ can be decomposed into 1-dimensional regional SD effect functions, as in Eq. (8) if all interaction-related heterogeneity is reduced. In this special case and if Shapley values are estimated by the marginal-based approach, it can be shown that $f^{SD,c}_{j|\mathcal{A}_{g}}=f^{PD,c}_{j|\mathcal{A}_{g}}$ (see Appendix C.3 for more details and an illustration in a simulated example).
@@ 4.6 Illustrative Comparison of Feature Effect Methods Within GADGET
Definitions and descriptions of the mean and interaction-related heterogeneity estimates and visualization techniques for each feature effect method can be found in Appendix C.4.
@@ 4.7 Quantifying Feature Interactions
$I(\mathcal{A}_{P},\tilde{\mathbf{x}}_{j})=\frac{\mathcal{R}_{j}(\mathcal{A}_{P},\tilde{\mathbf{x}}_{j})-\mathcal{R}_{j}(\mathcal{A}_{l},\tilde{\mathbf{x}}_{j})-\mathcal{R}_{j}(\mathcal{A}_{r},\tilde{\mathbf{x}}_{j})}{\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j})},$ (12)
@@
$I_{z}=\frac{\sum_{\mathcal{A}_{P}\in\mathcal{B}_{z}}\sum_{j\in S}(\mathcal{R}_{j}(\mathcal{A}_{P},\tilde{\mathbf{x}}_{j})-\mathcal{R}_{j}(\mathcal{A}_{l},\tilde{\mathbf{x}}_{j})-\mathcal{R}_{j}(\mathcal{A}_{r},\tilde{\mathbf{x}}_{j}))}{\sum_{j\in S}\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j})}.$ (14)
@@
$R^{2}_{j}=\sum_{z\in Z}I_{z,j}(\tilde{\mathbf{x}}_{j})=1-\frac{\sum_{\mathcal{A}_{t}\in\mathcal{B}_{t}}\mathcal{R}_{j}(\mathcal{A}_{t},\tilde{\mathbf{x}}_{j})}{\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j})},$ (15)
@@
$R^{2}_{Tot}=\sum_{z\in Z}I_{z}=1-\frac{\sum_{\mathcal{A}_{t}\in\mathcal{B}_{t}}\sum_{j\in S}\mathcal{R}_{j}(\mathcal{A}_{t},\tilde{\mathbf{x}}_{j})}{\sum_{j\in S}\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j})}.$ (16)
@@
5 A Procedure to Detect Global Feature Interactions
@@
As shown in Section 4.7, GADGET quantifies feature interactions between the feature subsets $S$ and $Z$. Choosing the features to include in $S$ and $Z$ depends on the underlying research question. If the user is interested in how a specific set of features ($S$) influences the model’s predictions depending on another user-defined feature set ($Z$), then $S$ and $Z$ are chosen based on domain knowledge. However, the user may not know which relationships the ML model has learned. Thus, choosing $S$ and $Z$ based on domain knowledge does not ensure consideration of all interacting features. If our goal is to minimize feature interactions between all features to potentially additively decompose the prediction function into univariate effects, we can define $S=Z=\{1,\ldots,p\}$. With that choice, we aim to reduce the overall interaction-related heterogeneity in all features ($S$), since we also consider all features to be possible interacting features ($Z$). For high-dimensional settings, this has the disadvantage of being computationally expensive and we might “detect” spurious interactions and therefore produce less stable regional splits. Hence, we must define beforehand the subset of features that actually interact. Given that features typically interact with each other and all involved features will usually show heterogeneity in their local effects while being responsible for the heterogeneity of other involved features, we will choose $S=Z$. Thus, we will only use $S$ as the globally interacting feature subset to be defined.
5.1 The PINT Procedure
We introduce a new permutation-based interaction detection (PINT) procedure to define the interacting feature subset $S$. PINT is based on the idea of a non-parametric permutation test (Ernst, 2004). It can be applied to any feature effect method satisfying Axiom 1.
With the risk function defined in Eq. (7) and based on the chosen local feature effect method, we can quantify the interaction-related heterogeneity of each feature $\mathbf{x}_{j}$ within the feature space $\mathcal{X}$. Due to correlations between features, the estimated heterogeneity might also include spurious interactions. Thus, we must define a null distribution to determine which heterogeneity is actually due to feature interactions (i.e., significant w.r.t. the null distribution) and which heterogeneity is due to other reasons, such as correlations or noise. Thus, the null distribution needs to retain the underlying properties of the data.
@@ 5.1 The PINT Procedure
Generally, we do not know the true functional relationship and there are infinitely many distributions compliant with a null hypothesis of no interaction-related heterogeneity: $\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j})=0$. We use the prediction model and a permutation-based approach to approximate the variability of the risk $\mathcal{R}_{j}$ under this null hypothesis. We follow a similar approach to Altmann et al. (2010) who suggested a permutation-based null distribution to detect important features. To obtain the null distribution, the target variable $y$ is permuted (line 4 in Algorithm 2), which breaks the association between the target and the features while keeping the feature distribution intact. Hence, no more interactions are present between $\mathbf{x}_{j}$ and $\mathbf{x}_{-j}$, but the association between the features remains. Thus, when performing a refit on the data set $\tilde{\mathcal{D}}$ with the permuted target variable, potential heterogeneity of local effects is only due to randomly learned effects of the prediction model depending on the feature distribution and noise and not caused by true feature interactions. Therefore, calculating ~$\tilde{\mathcal{R}}_{j}$ based on $\tilde{\mathcal{D}}$ measures the risk of $\mathbf{x}_{j}$ under the null (lines 5-8 in Algorithm 2). To obtain the null distribution of risk values for $\mathbf{x}_{j}$, this procedure is repeated $s$ times. [^8] An alternative approach would be to randomly permute $\mathbf{x}_{j}$ instead of $y$ to break the interaction effects between $\mathbf{x}_{j}$ and $\mathbf{x}_{-j}$. However, this approach is computationally more expensive and does not only break the respective interaction effects but also the correlations between $\mathbf{x}_{j}$ and $\mathbf{x}_{-j}$, and thus spurious interactions will not be detected. Using conditional samples (Molnar et al., 2023) or knock-offs (Watson and Wright, 2021) instead of random permutations solves the second problem. However, this approach is even more computationally expensive and becomes infeasible with a higher number of features.
@@
Since we are interested in defining the feature subset $S$ of all interacting features, we calculate ~$\tilde{\mathcal{R}}_{j}$ for all features $j\in\{1,\ldots,p\}$. Since all feature interactions between $y$ and $\mathbf{x}$ are broken by permuting $y$, the null distribution of risk values can be estimated for all features by the described $s$ permuting and refitting steps. To determine the interacting feature subset $S$, the respective risk value $\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}})$, is calculated based on the unpermuted data set $\mathcal{D}$ and the originally fitted model on this data set. The value of $\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}})$ is then compared to the $(1-\alpha)$-quantile, i.e., $z^{1-\alpha}_{j}$, of the non-parametric null distribution. If the risk value of $\mathbf{x}_{j}$ exceeds the respective quantile, then feature interactions between $\mathbf{x}_{j}$ and $\mathbf{x}_{-j}$ are considered to be highly relevant (lines 11-13 in Algorithm 2).
@@
risk function $\mathcal{R}_{j}$, significance level $\alpha$
@@
5:     refit model on $\tilde{\mathcal{D}}^{k}$ to obtain the prediction function $\tilde{f}^{k}_{\tilde{\mathcal{D}}^{k}}$
@@
7:        calculate risk ~$\tilde{\mathcal{R}}_{j}^{k}=\mathcal{R}_{j}^{k}(\mathcal{X},\tilde{\mathbf{x}}_{j},\tilde{f}^{k}_{\tilde{\mathcal{D}}^{k}})$ for the $j$-the feature achieved by $\tilde{f}^{k}_{\tilde{\mathcal{D}}^{k}}$
@@
11:     calculate risk $\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}})$ for $j$-the feature based on $\mathcal{D}$ and $\hat{f}$
@@
13:     add $j$ to $S$, if and only if $\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}})>z^{1-\alpha}_{j}$
@@
5.2 Theoretical Properties
We emphasize that PINT is a detection procedure, not a formal statistical test. This distinction is necessary because the permutation approach generates data from only one of the many possible distributions compliant with the null hypothesis. We may nevertheless ask how the procedure’s type I and type II errors behave. In the following, we give conditions under which PINT is valid as a statistical test. An empirical validation is given in Section 7 and Appendix F.
A difficulty in formally analyzing the procedure is that it depends on the unspecified ML algorithm that induces $\hat{f}$. We can emphasize this by writing $\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j},\hat{f}_{\mathcal{D}_{n}})$ for the risk for feature value $\mathbf{x}_{j}$ achieved by the algorithm $\hat{f}$ run on a data set [^9] With a slight abuse of notation, we use $\mathcal{D}_{n}$ to define the data set as a random variable. $\mathcal{D}_{n}=\{(X^{(i)},Y^{(i)})\}_{i=1}^{n}$, generated i.i.d. from $\mathds{P}_{X,Y}$. Define $f(\mathbf{x})=\mathds{E}[Y\mid X=\mathbf{x}]$ and $\mathcal{D}^{\Pi}_{n}=\{(X^{(i)},Y_{\Pi(i)}^{(i)})\}_{i=1}^{n}$ where $\Pi$ is a random permutation of $\{1,\dots,n\}$ independent of $\mathcal{D}_{n}$. We can now impose abstract, but interpretable conditions on the ML algorithm under which the test is valid.